# Deep learning-based species-area models reveal multi-scale patterns of species richness and turnover

### v1

@@ 2.3 Scale-specific patterns of plant species richness and turnover in Europe
Projections of species richness and turnover across spatial scales are essential for comparing biodiversity trends from local to regional levels. Given the superior performance of the deep SAR model informed with location and area-specific environmental features, we used it to map species richness and turnover across Europe at multiple spatial scales. We defined turnover as the slope of the SAR predicted at each location ($\frac{dS}{dA}$). Using a moving window approach with different spatial grains, we derived environmental features that served as predictors to estimate species richness at the centroid of each window. We also estimated location-specific SARs for three environmentally distinct locations by incrementally increasing the area of spatial units centered at each location of interest.
The model revealed markedly different patterns of species rihness across spatial scales. At fine spatial scales (1 km), mountainous areas exhibited relatively lower species richness compared to lowland regions (Fig. 4b), contrary to established biogeographical knowledge of European vascular plant diversity (Cai et al., 2023). However, predictions at coarser spatial scales (50 km) revealed the expected highest species richness along the Mediterranean coastline and in mountainous regions including the Alps, Carpathians, and Dinarides (Fig. 4c). This scale-dependent variation underscores the critical importance of considering spatial grain in macroecological analyses (Wiens, 1989; Portier et al., 2022).
The predicted location-specific SARs explicitly encapsulated these scale-dependent dynamics (Fig. 4a). The Southwestern Alps (blue curve) exhibited lower local species richness at fine spatial scales compared to the Northern Iberian Peninsula (purple curve) and Northern Germany (red curve). However, as spatial scale increases, species richness in the Alps surpasses both other locations due to higher species turnover. This pattern reflects the characteristic biodiversity structure of mountain environments, where lower local carrying capacity is offset by higher niche diversity driven by environmental heterogeneity, resulting in low local richness but high species turnover and regional richness (Antonelli et al., 2018). The geographic variations in species turnover reflected these patterns (Fig. 4d-e). Notably, turnover patterns differed substantially from species richness patterns at fine scales. Mountain environments showed moderate species richness but high turnover, which effectively translated into higher species richness at coarser scales.
@@ 3 Discussion
Our deep SAR model successfully predicted species rarefaction curves across spatial units of varying locations and areas by leveraging environmental features, allowing to predict location-specific species area relationships under asymptotic sampling effort. By predicting species rarefaction curves in environmental space, our approach leverages well-sampled regions to make predictions in environmentally similar but less-surveyed areas. This represents a significant advantage over conventional non-parametric estimators that rely solely on local incidence data without incorporating environmental context. Consequently, our model achieved a 32% improvement over the Chao2 estimator when predicting total species richness. Environmental features proved critical for model performance, reducing prediction error by 43% compared to area-only models. Shapley value analysis revealed that environmental conditions are the primary drivers of species richness across all spatial scales, while highlighting the increasing relative importance of area effects and environmental heterogeneity at larger spatial grains. Spatially explicit predictions of species richness and turnover uncovered scale-dependent biodiversity patterns, with Mediterranean coastlines and mountainous regions exhibiting highest species richness at coarse scales, despite mountainous areas showing relatively lower local richness at fine scales. These results collectively demonstrate that our approach can effectively capture the multi-scale relationships between environmental drivers and species richness patterns.
@@
The application of machine learning in quantitative ecology has predominantly focused on predicting species distributions (Brun et al., 2024). However, many ecological and conservation questions require understanding macroecological properties at the community or ecosystem level. Our deep SAR model adopts a fundamentally different approach by directly modeling community-level biodiversity patterns rather than individual species distributions. By capturing how environmental conditions influence species richness accumulation across spatial scales without explicitly modeling individual species’ ranges, our approach offers several advantages. First, it is more data-efficient as it leverages the aggregated signal from multiple species rather than requiring sufficient data for each species individually. Second, it can exploit general relationships between environmental heterogeneity and species richness that remain consistent across biogeographic regions (Leroux et al., 2017; Dubuis et al., 2011). This enhances model transferability and generalization compared to species distribution models, which often suffer from limited occurrence data, particularly for rare species, and may not capture emergent community-level properties that arise from species interactions and assembly processes.
@@
The theoretical functional form of the SAR has long been debated in ecology, with no consensus on the most appropriate model (Dengler, 2009). While the search for appropriate functional forms is crucial in low-data regimes where strong inductive bias is required, theoretical forms are simplifications that cannot accommodate the complexity of realistic landscapes. Within local species communities, species richness is expected to increase smoothly as sampling area grows, reflecting saturation of local species pools (Connor and McCoy, 1979; Hubbell, 2001). However, as sampling area expands to include neighboring distinct communities, species numbers may rise sharply due to incorporation of novel species assemblages (Whittaker, 1972; Gering et al., 2003). By leveraging abundant small-scale ecological survey data, we can employ more flexible deep neural networks to capture these complex dynamics without imposing strong assumptions about functional form. These models have the capacity to express nonlinear dynamics occurring across dispersal barriers and in response to complex natural or anthropogenic drivers across the full range of ecologically relevant scales.
@@
Our deep SAR model provides a comprehensive framework for biodiversity assessment and conservation planning by quantitatively evaluating environmental driver effects on species richness across spatial scales. Four instance, our deep SAR model could enable habitat loss impact assessment by comparing predicted species richness between intact and degraded areas (Gotelli and Colwell, 2001; Matthews et al., 2021a; Thomas et al., 2004; Brooks et al., 2002), while capturing scale-dependent effects often overlooked in traditional assessments. Species turnover maps offer particularly valuable conservation insights. High-turnover regions indicate rapid species richness increase with area, suggesting that habitat loss may disproportionately impact regional biodiversity by eliminating unique assemblages (Socolar et al., 2016). Conversely, areas with high local richness but low turnover may exhibit greater resilience to localized habitat loss, harboring species pools less dependent on large-scale environmental gradients. Explainable AI methods, such as Shapley values, quantify how environmental stressor effects on species richness vary across spatial scales, therefore allowing to enhance our understanding of biodiversity responses. The multi-scale capability of our model delivers critical information that can match the spatial scale of conservation objectives (Kunin et al., 2018).
@@
From a machine learning perspective, predicting species richness from pools of ecological surveys represents a multiple instance learning (MIL) problem (Carbonneau et al., 2018). Ecological surveys correspond to instances, their aggregation generates training samples (bags), and predicting species richness involves predicting bag-level labels from instance-specific (sampling effort) and bag-specific (environmental covariates) features. With this perspective, modern MIL architectures using attention mechanisms could potentially improve species richness estimates (Ilse et al., 2018). While we manually engineered environmental features by calculating means and standard deviations, convolutional neural networks or vision transformers could automatically extract higher-order statistics and leverage satellite imagery to characterize habitat fragmentation effects (Hanski et al., 2013; Rybicki and Hanski, 2013). Future improvements could leverage higher-resolution climate data to capture microenvironmental factors shaping local species assemblages (Dembicz et al., 2021), incorporate alternative biodiversity data sources such as citizen science species occurrence data (GBIF: The Global Biodiversity Information Facility, 2022), and extend predictions to other biodiversity metrics including Hill numbers that generalize species richness (Chao et al., 2014) or genetic diversity (De Kort et al., 2021). Enhanced uncertainty characterization beyond model ensembling represents another important direction for transparent biodiversity reporting.
@@ 4 Conclusion
Our study demonstrates how deep learning can effectively predict macroecological properties by leveraging abundant small-scale ecological survey data and sampling theory. The deep SAR model provides a data-efficient approach that delivers unprecedented insights into biodiversity patterns across spatial scales, revealing how environmental drivers operate differently at local versus regional scales. The model’s capacity to capture multi-scale biodiversity dynamics is essential for attributing biodiversity changes to global change drivers and designing comprehensive conservation strategies that account for the spatial scale-dependency of ecological processes.
@@ 5 Methods
5.1 Problem statement
Estimating total species richness from small-scale ecological surveys.
We aim to predict the total species richness $S_{\text{tot}}$ within a spatial unit of arbitrary location and area. Let $\lambda_{j}>0$ denote the density (individuals per unit area) of species $j\in\{1,\dots,S_{\text{tot}}\}$. The probability of observing $x_{j}$ individuals of species $j$ in an ecological survey of area $a$ follows a Poisson distribution, $P(N_{j}=x_{j}\mid a)=\frac{(\lambda_{j}a)^{x_{j}}e^{-\lambda_{j}a}}{x_{j}!}$. For $M$ randomly distributed ecological surveys with areas $\{a_{i}\}_{i=1}^{M}$, the probability of not detecting species $j$ in survey $i$ is $P(N_{j}=0\mid a_{i})=e^{-\lambda_{j}a_{i}}$. Assuming independence across surveys, the probability of detecting species $j$ in at least one of $m\leq M$ randomly selected surveys is
$1-\prod_{i=1}^{m}e^{-\lambda_{j}a_{i}}=1-e^{-\lambda_{j}\sum_{i=1}^{m}a_{i}},$ (1)
where $\sum_{i=1}^{m}a_{i}$ is the cumulative area of a specific realization of $m$ surveys. Considering the expected cumulative area when sampling $m$ surveys uniformly at random, denoted by $A(m)=\mathbb{E}\left[\sum_{i=1}^{m}a_{i}\right]$, the expected observed species richness $\mathbb{E}[S_{\text{obs}}(m)]$ in $m$ randomly selected surveys is:
$\mathbb{E}[S_{\text{obs}}(m)]=\sum_{j=1}^{S_{\text{tot}}}\left[1-e^{-\lambda_{j}A(m)}\right].$ (2)
[…]
Maximum likelihood estimation of the rarefaction curve.
In practice, we do not know the distribution $f$ nor the total species richness $S_{\text{tot}}$. Instead, parametric approaches assume the existence of a functional form $g_{\Theta}(m)$ that is monotonically increasing and exhibits asymptotic behavior, approximating the expected species richness $\mathbb{E}[S_{\text{obs}}(m)]$ for a given number $m$ of randomly selected surveys (SoberónM. and LlorenteB, 1993; Gotelli and Colwell, 2001; Colwell et al., 1997), where $\Theta$ represents the model parameters. Estimating the asymptote reduces to finding $\Theta$ such that $g_{\Theta}(m)$ accurately reproduces the observed species richness $S_{\text{obs}}(m)$ across various values of $m$. Given observed species counts $\{S_{\text{obs}}^{(k)}(m_{k})\}_{k=1}^{K}$ from $K$ random subsets of surveys with sizes $m_{1},\dots,m_{K}$, we can estimate $\Theta$ through maximum likelihood estimation. Notably, the choice of model $g_{\Theta}(m)$ implicitly determines an assumption about the distribution of species densities $f$ For instance, the commonly used Michaelis-Menten model can be derived from Eq. 3 by assuming that species densities follow an exponential distribution (see Derivation of the Michaelis-Menten model from the exponential distribution.).

### v2

@@ 2.3 Scale-specific patterns of plant species richness and turnover in Europe
Projections of species richness and turnover across spatial scales are essential for comparing biodiversity trends from local to regional levels. Here we used our deep SAR model to map species richness and turnover across Europe at multiple spatial scales. We defined turnover as the slope of the SAR predicted at each location ($\frac{dS}{dA}(A)$). Using a moving window approach with different spatial grains, we derived environmental features that served as predictors to estimate species richness within each window. We also estimated location-specific SARs for three environmentally distinct locations by incrementally increasing the area of spatial units centered at each location of interest.
The model revealed markedly different patterns of species richness across spatial scales. At fine spatial scales (1 km), mountainous areas exhibited relatively lower species richness compared to lowland regions (Fig.˜4b), contrary to established biogeographical knowledge of European vascular plant diversity (Cai et al., 2023). However, predictions at coarser spatial scales (50 km) revealed the expected highest species richness along the Mediterranean coastline and in mountainous regions including the Alps, Carpathians, and Dinarides (Fig.˜4c). This scale-dependent variation underscores the critical importance of considering spatial grain in macroecological analyses (Wiens, 1989; Portier et al., 2022).
The predicted location-specific SARs explicitly embed these scale-dependent dynamics (Fig.˜4a). The Southwestern Alps (blue curve) exhibited lower local species richness at fine spatial scales compared to the Northern Iberian Peninsula (purple curve) and Northern Germany (red curve). However, as spatial scale increases, species richness in the Alps surpasses both other locations due to higher species turnover. This pattern reflects the characteristic biodiversity structure of mountain environments, where lower local carrying capacity is offset by higher niche diversity driven by environmental heterogeneity, resulting in low local richness but high species turnover and regional richness (Antonelli et al., 2018). The geographic variations in species turnover reflected these patterns (Fig.˜4d-e). Notably, turnover patterns differed substantially from species richness patterns at fine scales. Mountain environments showed moderate species richness but high turnover, which effectively translated into higher species richness at coarser scales.
@@ 3 Discussion
In this work, we demonstrated how deep learning combined with sampling theory enables spatially explicit prediction of species richness from fine to coarse spatial grains. Our deep SAR model learns species rarefaction curves conditioned on spatial unit area and associated environmental features, allowing to leverage well-sampled regions to make predictions in environmentally similar but less-surveyed areas. By extrapolating rarefaction curves under asymptotic sampling effort, our approach yields estimates of total expected species richness for any spatial unit area, thereby recovering location-specific species-area relationships. Incorporating environmental context contributed to improve species richness estimates by 32% over the conventional Chao2 estimator. Shapley value analysis of the deep SAR model revealed insights into key drivers of biodiversity patterns. The analysis supported that environmental filtering is the primary driver of species richness patterns, with area per se effects and habitat diversity effects becoming increasingly important at larger spatial grains. Spatially explicit predictions of species richness and turnover uncovered scale-dependent biodiversity patterns, with mountainous regions exhibiting the highest species richness at coarse scales, despite showing relatively lower local richness at fine scales. These results collectively demonstrate that our approach effectively captures the multi-scale relationships between biodiversity drivers and species richness patterns.
@@
The application of machine learning in macroecology research has predominantly focused on predicting species distributions (see, e.g., (Deneu et al., 2021; Brun et al., 2024; Gillespie et al., 2024)). However, many ecological and conservation questions require understanding macroecological properties at the community or ecosystem level. Our deep SAR model adopts a fundamentally different approach by directly modeling community-level biodiversity metrics rather than individual species distributions. By capturing how environmental conditions influence species richness across spatial scales without explicitly modeling individual species’ ranges, our approach offers several advantages. First, it is more data-efficient as it leverages the aggregated signal from multiple species rather than requiring sufficient data for each species individually. Second, it can exploit general relationships between environmental conditions and species richness that remain consistent across biogeographic regions (Leroux et al., 2017; Dubuis et al., 2011). This enhances model transferability and generalization compared to species distribution models, which often suffer from limited occurrence data, particularly for rare species, and may not capture emergent community-level properties that arise from species interactions and assembly processes.
@@
The theoretical functional form of the SAR has long been debated in ecology, with no consensus on the most appropriate model (Dengler, 2009). While the search for a theoretical functional forms is required in a low-data regime or for deriving general macroecological patterns, they are simplifications that cannot accommodate the complexity of realistic landscapes. Realistic landscapes are often heterogeneous, where species richness may rise idiosyncratically due to dispersal barriers, natural local variations, or anthropogenic stressors (Whittaker, 1972; Gering et al., 2003). Sampling theory allows us to leverage abundant ecological survey data for training deep neural networks, which have the capacity to capture these complex accumulation dynamics without imposing strong assumptions about the functional form.
@@
Our deep SAR model provides a comprehensive framework for biodiversity assessment and conservation planning by enabling the estimation of environmental driver effects on species richness across spatial scales. For instance, our deep SAR model could enable habitat loss impact assessment by comparing predicted species richness between intact and degraded areas (Gotelli and Colwell, 2001; Matthews et al., 2021a; Thomas et al., 2004; Brooks et al., 2002), while capturing scale-dependent effects often overlooked in traditional assessments. Additionally, species turnover maps offer particularly valuable conservation insights. High-turnover regions indicate rapid species richness increase with area, suggesting that habitat loss may disproportionately impact regional biodiversity by eliminating unique assemblages (Socolar et al., 2016). Conversely, areas with high local richness but low turnover may exhibit greater resilience to localized habitat loss. Explainable artificial intelligence methods, such as Shapley values, enable us to quantify how environmental stressor effects on species richness vary across spatial scales, therefore allowing to enhance our understanding of biodiversity responses. Overall, our deep SAR model delivers critical information that can serve the multi-scale objectives of biodiversity conservation (Kunin et al., 2018).
@@
From a machine learning perspective, predicting species richness from pools of ecological surveys represents a multiple instance learning (MIL) problem (Carbonneau et al., 2018). Ecological surveys correspond to instances, their aggregation generates training samples (bags), and predicting species richness involves predicting bag-level labels from instance-specific (sampling effort) and bag-specific (environmental covariates) features. With this perspective, modern MIL architectures using attention mechanisms could potentially improve species richness estimates (Ilse et al., 2018). While we manually engineered environmental features by calculating means and standard deviations, convolutional neural networks or vision transformers could automatically extract higher-order statistics and leverage satellite imagery to characterize habitat fragmentation effects (Hanski et al., 2013; Rybicki and Hanski, 2013). Future improvements could leverage higher-resolution climate data to capture microenvironmental factors shaping local species assemblages (Dembicz et al., 2021), incorporate alternative biodiversity data sources such as citizen science species occurrence data (GBIF: The Global Biodiversity Information Facility, 2022), and extend predictions to other biodiversity metrics including Hill numbers that generalize species richness (Chao et al., 2014) or genetic diversity (De Kort et al., 2021). Enhanced uncertainty characterization beyond model ensembling represents another important direction for transparent biodiversity assessment.
@@ 4 Conclusion
Our study demonstrates how deep learning combined with sampling theory can effectively estimate scale-dependent macroecological properties from small-scale yet abundant ecological data. The deep SAR model provides a data-efficient approach to deliver comprehensive insights into biodiversity patterns, revealing the distinct effects of biodiversity drivers across scales. We anticipate future applications of our framework to capture complementary macroecological properties of ecosystems, such as functional or genetic diversity. This multi-scale prediction capacity is essential for attributing biodiversity changes to global change drivers and designing robust conservation strategies that account for the scale-dependency of ecological processes.
@@ 5 Methods
We aim to estimate total species richness $S_{\mathrm{T}}$, the total number of species in a spatial unit $\mathsf{A}$} at a given location. However, biodiversity surveys typically sample only a small fraction of $\mathsf{A}$}, requiring extrapolation from these limited samples to estimate richness across the entire area. This extrapolation challenge has been addressed through various approaches (Elena Schmitz and Rahmann, 2025), but these methods typically focus on estimating total richness within a single spatial unit. Our approach introduces a key innovation by conditioning the extrapolation on features $X$ associated with the spatial unit, including its area and statistics of environmental features within the spatial unit. This conditioning on $X$ enables spatially explicit mapping of $S_{\mathrm{T}}=S_{\mathrm{T}}(X)$ estimates across landscapes at multiple spatial scales.
Section˜5.1 provides the required definitions and explains our methodological setup. Section˜5.2 details the curve fitting procedure and neural network implementation. Section˜5.3 describes the datasets used in our analysis.
5.1 Estimating total species richness from small samples
Sampling process.
Let $\mathsf{a}_{\boldsymbol{i}}$ be the union of a set of $m=|{\boldsymbol{i}}|$ small-scale samples $\mathsf{a}_{i_{1}},...,\mathsf{a}_{i_{m}}$ randomly placed in the spatial unit $\mathsf{A}$}:
$\mathsf{a}_{\boldsymbol{i}}:=a_{i_{1}}\cup\cdots\cup a_{i_{m}}\subseteq{\mathsf{A}},$ (1)
with corresponding locations $x_{i_{1}},...,x_{i_{m}}$ and pooled area $|\mathsf{a}_{\boldsymbol{i}}|$. We will denote such a set as a sample set and view a particular sample set as a realisation of a random variable. In the limit of infinite sample effort, we have
$\lim_{m\to\infty}\mathsf{a}_{\boldsymbol{i}}=\lim_{m\to\infty}(a_{i_{1}}\cup\cdots\cup a_{i_{m}})={\mathsf{A}}.$ (2)
In the sample set, we count species to obtain its observed species richness:
$S(\mathsf{a}_{\boldsymbol{i}}):=\sum_{j=1}^{S_{\mathrm{T}}}\mathds{1}_{N_{j}(\mathsf{a}_{\boldsymbol{i}})\geq 1},$ (3)
where $\mathds{1}_{(\cdot)}$ evaluates to 1 when the condition in its argument is true and to zero otherwise, and $N_{j}(\mathsf{a}_{\boldsymbol{i}})$ the number of individuals of species $j$ in $\mathsf{a}_{\boldsymbol{i}}$. Randomly sampling over different $\boldsymbol{i}$} yields pairs $|\mathsf{a}_{\boldsymbol{i}}|\mapsto S(\mathsf{a}_{\boldsymbol{i}})$. To avoid bias towards any particular $\boldsymbol{i}$}, we need a way to average over different realisations.
Species rarefaction curve.
The curve that shows how expected species richness increases with sampling effort is known as a rarefaction curve. We can obtain it by taking an expectation conditional on sampling effort, where sampling effort may either be expressed in terms of sample size $m=|{\boldsymbol{i}}|$ or in terms of pooled sample area $|\mathsf{a}_{\boldsymbol{i}}|$. Because we will fit a species-area relation, we use the latter:
$\mathcal{S}(a):=\mathbb{E}\left[S(\mathsf{a}_{\boldsymbol{i}})\Big{|}|\mathsf{a}_{\boldsymbol{i}}|=a\right]=\sum_{j=1}^{S_{\mathrm{T}}}\mathbb{P}\left[N_{j}(\mathsf{a}_{\boldsymbol{i}})\geq 1\Big{|}|\mathsf{a}_{\boldsymbol{i}}|=a\right],$ (4)
where we used Eq.˜3 and that the expectation of an indicator function is the probability of its argument. Defining
$p_{k}(a):=\frac{1}{S_{\mathrm{T}}}\sum_{j=1}^{S_{\mathrm{T}}}\mathbb{P}\left[N_{j}(\mathsf{a}_{\boldsymbol{i}})=k\Big{|}|\mathsf{a}_{\boldsymbol{i}}|=a\right],$ (5)
the expected proportion of species in a sample with $k$ individuals at effort $a$, we obtain
$\mathcal{S}(a)=S_{\mathrm{T}}\sum_{k=1}^{\infty}p_{k}(a)=S_{\mathrm{T}}[1-p_{0}(a)],$ (6)
where we used that Eq.˜5 sums to 1. As $S_{T}$ denotes the number of species occurring in $\mathsf{A}$}, $p_{0}(a)$ is the expected fraction of species occurring in $\mathsf{A}$} that are missed at effort $a$. We can write $p_{0}$ and therefore $\mathcal{S}(a)$ as (see SI Section˜S1.1; Harte et al. (2009))
$p_{0}(a)=\sum_{n\geq 1}f^{+}_{n}q_{n}(a),\quad\Rightarrow\quad\mathcal{S}(a)=S_{\mathrm{T}}[1-\sum_{n\geq 1}f^{+}_{n}q_{n}(a)].$ (7)
where $f^{+}_{n}$ is the zero-truncated species abundance distribution (SAD, over species present) in $\mathsf{A}$} and $q_{n}(a)$ the probability of absence at effort $a$ given $n$ individuals in $\mathsf{A}$}. $q_{n}$ depends only on how individuals are arranged within $\mathsf{A}$}. Thus the rarefaction curve depends jointly on the SAD, the spatial arrangement of individuals within $\mathsf{A}$} and the sampling process. By construction $\mathcal{S}(0)=0$ and $\mathcal{S}(|{\mathsf{A}}|)=S_{\mathrm{T}}$.
Parametric estimation of the rarefaction curve.
We aim to estimate $\mathcal{S}(a)$ by fitting functional forms to sampling data. In our choice of functional form, the case of uniform random placement (URP) is informative because it yields closed form solutions for the rarefaction curve in terms of $a$. Under URP, we have $q_{n}(a)=(1-a/|{\mathsf{A}}|)^{n}$, such that the rarefaction curve becomes (via Eq.˜7; Coleman (1981); He and Legendre (2002)):
$\mathcal{S}_{\mathrm{u}}(\alpha;\vartheta)=S_{\mathrm{T}}\bigl{[}1-G\{f^{+}_{n,\vartheta}\}(1-\alpha)\bigr{]},$ (8)
where $\alpha:=a/|{\mathsf{A}}|$ and $G\{f^{+}_{n,\vartheta}\}$ the probability-generating function (PGF) of the SAD with parameters $\vartheta$. Table˜S1 shows examples for a selection of SADs under URP. For instance, when $f^{+}_{n,\vartheta}$ is a geometric distribution with parameter $\vartheta$, then the rarefaction curve is the classic Michaelis–Menten type functional form $\mathcal{S}_{\mathrm{u}}(\alpha;\vartheta)=S_{\mathrm{T}}\tfrac{\alpha}{\vartheta+(1-\vartheta)\alpha}$. The URP case suggests that the functional form of the rarefaction equation can be obtained by first choosing the right parametric family of probability mass functions $f^{+}_{n,\vartheta}$ with parameter(s) $\vartheta$, then obtaining the corresponding form of the rarefaction curve via Eq.˜8, and finally obtaining the parameters $\Theta=(S_{\mathrm{T}},\vartheta)$ by fitting it to data generated by the sampling process.
Real communities, however, exhibit spatial structure due to environmental heterogeneity and spatial interactions, so URP is typically violated. Moreover both spatial structure and the SAD generally differ between locations, so it would be highly restrictive to choose a single parametric family of probability mass functions. Hence, any chosen functional form $g_{\Theta}(a)$ must be flexible enough as to fit diverse shapes regardless of the deviation from URP and differences between locations, but still satisfy monotonicity in $a$ and reach $S_{\mathrm{T}}$ for large $a$ ($m\to\infty$). Parameters $\Theta$ can be obtained by fitting $g_{\Theta}$ to data pairs $|\mathsf{a}_{\boldsymbol{i}}|\mapsto S(\mathsf{a}_{\boldsymbol{i}})$, generated through the sampling process described above, drawing $\mathsf{a}_{\boldsymbol{i}}$ randomly with replacement. Fitting can be performed using maximum likelihood estimation. Once $g_{\Theta}$ is fitted, we can recover $S_{T}$ as the limit of the rarefaction curve to large sampling effort.